# Predictive Pattern Recognition Techniques Towards Spatiotemporal Representation of Plant Growth in Simulated and Controlled Environments: A Comprehensive Review

### v2

@@ 5.3.1 Anomaly Detection
Anomaly detection and early interventions practices during crop development benefit from predictive pattern recognitions techniques that incorporate dynamic feedback from environmental growing conditions. The focus on the development of models that capture long-term temporal dependencies through sequence-based methods, such as transformers and RNNs, provides an effective approach to identifying deviations in plant growth trajectories. For instance, temporal anomalies such as stunted growth or disease occurrence can be detected in advance by analyzing sequential data from real-time HTPP platform measurements. Moreover, the incorporation multi-modal approaches to combine imaging data with contextual environmental parameters presents a more comprehensive assessment of plant health. These models improve the precision of anomaly detection and allow for dynamic adjustments through intervention and control strategies. Such techniques are particularly useful for mitigating potential yield losses and optimizing resource use in controlled environments and field systems.
@@ 5.3.2 Plant-Environment Simulation
Simulation frameworks informed by generative and probabilistic models improve the ability to predict plant growth under varying environmental conditions. Models like the CGMs presented in Section 4.2.2 enable the synthesis of realistic growth trajectories influenced by the spatial and dynamic growing condition, including the distribution of light, nutrient conditions, temperature, and CO2. This provides in silico methods to inform the structural development process of a controlled plant system and to develop strategies prior to conducting in-field operations. These frameworks allow for resource optimization in controlled environments and improve the forecasting of long-term plant responses to changing environments.
@@ 5.3.3 Digital Twinning and Representation
Digital twinning within plant phenomics typically consists of methods for 3D reconstruction of plant structures taken at sparse static snapshots within its growth cycle. This is often due to limitations in imaging techniques for capturing frequent 3D spatial data during its growth. The structured representation methods explored within this review have implications on expanding this static framework to capture a continuous virtual replica of a plant system during its entire growth cycle. The continuous frame synthesis techniques explored in Debbagh et al. [Debbagh2024SIPGS], allow for the interpolation of structured outputs between points of time. This presents an opportunity to generate digital plant structures with the guidance of few snapshots of plants. Moreover, the divergence of these plant structures from the snapshots can be explored through simulated growing conditions.

### v3

@@ 5.3.1 Anomaly Detection
Anomaly detection and early interventions practices during crop development benefit from predictive pattern recognitions techniques that incorporate dynamic feedback from environmental growing conditions. The focus on the development of models that capture long-term temporal dependencies through sequence-based methods, such as transformers and RNNs, provides an effective approach to identifying deviations in plant growth trajectories. For instance, temporal anomalies such as stunted growth or disease occurrence can be detected in advance by analyzing sequential data from real-time HTPP platform measurements [152]. Moreover, the incorporation of multi-modal approaches to combine imaging data with contextual environmental parameters presents a more comprehensive assessment of plant health. Models such as multi-variate transformers have been employed to improve the precision of anomaly detection in time-series data [153, 154]. Such techniques are particularly useful for early intervention, mitigating potential yield losses and optimizing resource use.
@@ 5.3.2 Plant-Environment Simulation
Simulation frameworks informed by generative and probabilistic models improve the ability to predict plant growth under varying environmental conditions. Models like the CGMs presented in Section 4.2.2 enable the synthesis of realistic growth trajectories influenced by the spatial and dynamic growing condition, including the distribution of light, nutrient conditions, temperature, and CO2 [137, 134]. This provides in silico methods to inform the structural development process of a controlled plant system and to develop strategies prior to conducting in-field operations. These frameworks allow for resource optimization in controlled environments and improve the forecasting of long-term plant responses to changing environments.
@@ 5.3.3 Digital Twinning and Representation
Digital twinning is the virtual replication of a physical system, maintained through continuous data exchange, simulation, and feedback. Within plant phenomics, digital twinning is limited and typically consists of methods for 3D reconstruction of plant structures taken at sparse static snapshots within its growth cycle. This is often due to limitations in imaging techniques for capturing frequent 3D spatial data during its growth and challenges in continuous collection of auxiliary data [142, 155]. The structured representation methods explored within this review have implications on expanding this static framework to capture a continuous virtual replica of a plant system during its entire growth cycle. The continuous frame synthesis techniques explored in Debbagh et al. [137], allow for the interpolation of structured outputs between points of time. Moreover, FSPMs such as L-systems also demonstrate a potential for continuous state estimation by basing structural development on biological rules as a method that does not rely on continuous data collection [156]. These methods present an opportunity to generate digital plant structures with the guidance of few snapshots of plants. Moreover, the divergence of these plant structures from the snapshots can be explored through simulated growing conditions.